# arXiv PreprintSubmitted to arXiv on February 6, 2025

### v2

@@ Abstract
The Ensemble Kalman Filter (EnKF) is a widely used method for data assimilation in high-dimensional systems. In this paper, we show that the ensemble update step of the EnKF is equivalent to an empirical version of the Matheron update popular in the study of Gaussian process regression.
While this connection is simple, it seems not to be widely known, the literature about each technique seems distinct, and connections between the methods are not exploited. This paper exists to provide an informal introduction to the connection, with the necessary definitions so that it is intelligible to as broad an audience as possible.
@@ 1 Introduction
The Ensemble Kalman Filter (EnKF) (Evensen, 2003, 2009) is a cornerstone in data assimilation for large-scale dynamical systems due to its computational efficiency and scalability. The EnKF approximates the state estimation problem by evolving an ensemble of state vectors through the model dynamics and updating them using observational data.
Separately, the Matheron update provides a sample-based method for conditioning Gaussian random variables on observations (Doucet, 2010; Wilson et al., 2020, 2021). This approach is well-established in geostatistics and spatial statistics but the connection to ensemble methods in data assimilation seems not to be well-known.
In this work, we establish that the ensemble update step in the EnKF is equivalent to an empirical Matheron update, by putting them on a common probabilistic footing. By explicitly representing the ensemble mean and covariance using empirical approximations, we demonstrate this equivalence. Recognizing this connection provides an alternative probabilistic foundation for the EnKF and suggests potential improvements in ensemble data assimilation techniques by leveraging the properties of the Matheron update. Conversely, the analytic Matheron updates in the literature could benefit from the many computational optimisations arising from the data assimilation community.
@@ 1.1 Notation
We write random variates sans serif, $\boldsymbol{\mathsf{x}}$. Equality in distribution is =$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}$. The law, or measure, of a random variate $\boldsymbol{\mathsf{x}}$ is denoted $\mu_{\boldsymbol{\mathsf{x}}}$, so that $\left(\mu_{\boldsymbol{\mathsf{x}}}=\mu_{\boldsymbol{\mathsf{y}}}\right)\Rightarrow\left(\boldsymbol{\mathsf{x}}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\boldsymbol{\mathsf{y}}\right)$. Mnemonically, samples drawn from the $\mu_{\boldsymbol{\mathsf{x}}}$ are written with a serif $\boldsymbol{x}\sim\mu_{\boldsymbol{\mathsf{[}}x]}$. We use a hat to denote empirical estimates, e.g. $\widehat{\mu}_{\mathrm{X}}$ is the empirical law induced by the sample matrix $\mathrm{X}$. When there is no ambiguity we suppress the sample matrix, writing simply $\widehat{\mu}$. [^1] This measure notation from probability theory is unpopular in both machine learning and data assimilation contexts where densities are usually implied. We include a translation in Appendix A.
@@ 2 Matheron Update
$\begin{bmatrix}\boldsymbol{\mathsf{x}}\\ \boldsymbol{\mathsf{y}}\end{bmatrix}$ $\sim\mathcal{N}\left(\begin{bmatrix}\boldsymbol{m}_{\boldsymbol{\mathsf{x}}}\\ \boldsymbol{m}_{\boldsymbol{\mathsf{y}}}\end{bmatrix},\begin{bmatrix}\mathrm{C}_{\boldsymbol{\mathsf{xx}}}&\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\\ \mathrm{C}_{\boldsymbol{\mathsf{yx}}}&\mathrm{C}_{\boldsymbol{\mathsf{yy}}}\end{bmatrix}\right),$ (1)
@@
$\left(\boldsymbol{\mathsf{x}}\mid\boldsymbol{\mathsf{y}}{=}\boldsymbol{y}^{*}\right)$ =$\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\boldsymbol{\mathsf{x}}+\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}\left(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}}\right).$ (2)
@@
$\operatorname{Var}\left[\boldsymbol{\mathsf{x}}+\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}})\right]$ $=\operatorname{Var}[\boldsymbol{\mathsf{x}}]+\operatorname{Var}[\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}})]$ (10)
$\hskip 20.00003pt+\operatorname{Var}(\boldsymbol{\mathsf{x}},\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}}))+\operatorname{Var}(\boldsymbol{\mathsf{x}},\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}}))^{\top}$ (11)
@@
This pathwise approach to conditioning Gaussian variates has gained currency in machine learning as a tool for sampling and inference in Gaussian processes (Wilson et al., 2020, 2021), notably in generalising to challenging domains such as Riemannian manifolds (Borovitskiy et al., 2020).
@@ 3 Kalman Filter
We assume a known observation operator $\mathcal{H}$ such that observations $\boldsymbol{\mathsf{y}}$ are a priori related to state $\boldsymbol{\mathsf{x}}$ by $\boldsymbol{\mathsf{y}}=\mathcal{H}(\boldsymbol{\mathsf{x}}\boldsymbol{\mathsf{)}}$; Thus there exists a joint law for the prior random vector
@@
$\begin{bmatrix}\boldsymbol{\mathsf{x}}\\ \boldsymbol{\mathsf{y}}\end{bmatrix}$ $\sim\mathcal{N}\!\left(\begin{bmatrix}\boldsymbol{m}_{\boldsymbol{\mathsf{x}}}\\ \boldsymbol{m}_{\boldsymbol{\mathsf{y}}}\end{bmatrix},\begin{bmatrix}\mathrm{C}_{\boldsymbol{\mathsf{xx}}}&\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\\ \mathrm{C}_{\boldsymbol{\mathsf{yx}}}&\mathrm{C}_{\boldsymbol{\mathsf{yy}}}\end{bmatrix}\right)$ (15)
$=\mathcal{N}\!\left(\begin{bmatrix}\boldsymbol{m}_{\boldsymbol{\mathsf{x}}}\\ \mathrm{H}\,\boldsymbol{m}_{\boldsymbol{\mathsf{x}}}\end{bmatrix},\begin{bmatrix}\mathrm{C}_{\boldsymbol{\mathsf{xx}}}&\mathrm{C}_{\boldsymbol{\mathsf{xx}}}\,\mathrm{H}^{\top}\\ \mathrm{H}\,\mathrm{C}_{\boldsymbol{\mathsf{xx}}}&\mathrm{H}\,\mathrm{C}_{\boldsymbol{\mathsf{xx}}}\,\mathrm{H}^{\top}+\mathrm{R}\end{bmatrix}\right)$ (16)
@@
$\mathrm{K}$ :-$\coloneq\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\,\mathrm{H}^{\top}\left(\mathrm{C}_{\boldsymbol{\mathsf{yy}}}\right)^{-1},:$ (22)
:-$\coloneq\mathrm{C}_{\boldsymbol{\mathsf{xx}}}\,\mathrm{H}^{\top}\left(\mathrm{H}\,\mathrm{C}_{\boldsymbol{\mathsf{xx}}}\,\mathrm{H}^{\top}+\mathrm{R}\right)^{-1},:$ (23)
is the Kalman gain. Herafter we consider a constant diagonal $\mathrm{R}$ for simplicity, so that $\mathrm{R}=\rho^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$.
@@ 4 Ensemble Kalman Filter
[…]
@@
$\widehat{\operatorname{Var}}(\boldsymbol{\mathsf{x}})\coloneq\operatorname{Var}_{\boldsymbol{\mathsf{x}}\sim\widehat{\mu}_{\mathrm{X}^{\prime}}}(\boldsymbol{\mathsf{x}})=\widehat{\mathrm{C}}_{\boldsymbol{\mathsf{xx}}}$ :-$\coloneq\frac{1}{N-1}\sum_{i=1}^{N}\left(\boldsymbol{x}^{(i)}-\widehat{\boldsymbol{m}}_{\boldsymbol{\mathsf{x}}}\right)\left(\boldsymbol{x}^{(i)}-\widehat{\boldsymbol{m}}_{\boldsymbol{\mathsf{x}}}\right)^{\top}:$ (27)
@@
This association should be understood as a mnemonic device; rigorous convergence of such empirical measures to the true distribution is beyond the scope of this short note, and is rather subtle (Le Gland et al., 2009; Mandel et al., 2011; Kelly et al., 2014; Kwiatkowski and Mandel, 2015; Del Moral et al., 2017). For now, we regard posterior inference about $\boldsymbol{\mathsf{x}}$ given $\boldsymbol{\mathsf{y}}=\boldsymbol{y}^{*}$ as successful if we can find a new ensemble $\mathrm{X}^{\prime}$ such that its empirical moments approximate the desired target distribution, i.e. we seek to obtain an ensemble $\mathrm{X}^{\prime}$ such that
@@
$\mathrm{Y}\coloneq\mathcal{H}\mathrm{X}$ :-$\coloneq\begin{bmatrix}\mathcal{H}(\boldsymbol{x}^{(1)})&\mathcal{H}(\boldsymbol{x}^{(2)})&\dots&\mathcal{H}(\boldsymbol{x}^{(N)})\end{bmatrix}.:$ (32)
@@
$\widehat{\mu}_{\left[\begin{smallmatrix}\mathrm{X}\\ \mathrm{Y}\end{smallmatrix}\right]}$ :-$\coloneq\mathcal{N}\left(\begin{bmatrix}\bar{\mathrm{X}}\\ \bar{\mathrm{Y}}\end{bmatrix},\begin{bmatrix}\breve{\mathrm{X}}\breve{\mathrm{X}}^{\top}+\xi^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{x}}}}&\breve{\mathrm{X}}\breve{\mathrm{Y}}^{\top}\\ \breve{\mathrm{Y}}\breve{\mathrm{X}}^{\top}&\breve{\mathrm{Y}}\breve{\mathrm{Y}}^{\top}+\upsilon^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}\end{bmatrix}\right).:$ (33)
@@
where $\rho^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$ is the observation error covariance matrix. The term $\upsilon^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$ comes from the regularization of the empirical observation covariance (as defined by the ensemble of $\mathrm{Y}$), while $\rho^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$ represents the known covariance of the observation noise. Combining these two gives the effective covariance in the observation space.
Note that although we are still treating the ensemble as a sample from a Gaussian distribution, the predictive will in general not be Gaussian if the observation operator is nonlinear. Nonetheless, since the moments are by definition well-defined for any empirical distribution, the ensemble update is still well-defined in this case. Practitioners in the field gloss over this point, and the EnKF is widely used in nonlinear settings based on empirical effectivenes. Practical convergence guarantees in nonlinearity seem ellusive.
@@
5 Empirical Matheron Update is equivalent to Ensemble Kalman Update
@@ 6 Computational Complexity
As is well understood in the EnKF literature, part of the utility of the EnKF lies in the fact that the analysis update can be performed using only empirical ensemble statistics, thereby avoiding the direct computation and inversion of high-dimensional covariance matrices. In contrast, the naive Kalman update must operate directly on these full covariance matrices, which can be prohibitively expensive in high-dimensional settings. This is unneeded in the EnKF; in practice we calculate the gain (34) as
$\widehat{\mathrm{K}}$ $=\widehat{\mathrm{C}}_{\boldsymbol{\mathsf{xx}}}\,\mathcal{H}^{\top}\left(\mathcal{H}\,\widehat{\mathrm{C}}_{\boldsymbol{\mathsf{xx}}}\,\mathcal{H}^{\top}+\gamma^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}\right)^{-1}$ (39)
$=\breve{\mathrm{X}}\breve{\mathrm{Y}}^{\top}\left(\breve{\mathrm{Y}}\breve{\mathrm{Y}}^{\top}+\gamma^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}\right)^{-1}$ (40)
$=\gamma^{-2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}-\gamma^{-4}\breve{\mathrm{Y}}\left(\mathrm{I}_{N}+\gamma^{-2}\breve{\mathrm{Y}}^{\top}\breve{\mathrm{Y}}\right)^{-1}\breve{\mathrm{Y}}^{\top}$ (41)
where $\gamma^{2}=\upsilon^{2}+\rho^{2}$ aggregates the regularization term for the empirical observation covariance ($\upsilon^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$) and the intrinsic observation noise ($\rho^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$). The overall computational cost of the ensemble update is thus
$\mathcal{O}\left(D_{\boldsymbol{\mathsf{y}}}N^{2}+N^{3}+D_{\boldsymbol{\mathsf{x}}}D_{\boldsymbol{\mathsf{y}}}N\right).$
Since in practical applications the ensemble size $N$ is typically much smaller than the dimensions $D_{\boldsymbol{\mathsf{x}}}$ and $D_{\boldsymbol{\mathsf{y}}}$, these costs are significantly lower than those incurred by the naive Kalman update which is dominated by $\mathcal{O}(D_{\boldsymbol{\mathsf{y}}}^{3})$ operations. By carefully ordering operations to work with the $N\times N$ system, we thus achieve an update that is both computationally efficient and scalable to high-dimensional state spaces.

### v3

# arXiv PreprintSubmitted to arXiv on August 8, 2025

@@ Abstract
The Ensemble Kalman Filter (EnKF) is a widely used method for data assimilation in high-dimensional systems, with an ensemble update step equivalent to an empirical version of the Matheron update popular in Gaussian process regression—a connection that links half a century of data-assimilation engineering to modern path-wise GP sampling.
This paper provides a compact introduction to this simple but under-exploited connection, with necessary definitions accessible to all fields involved.
Source code is available at https://github.com/danmackinlay/paper_matheron_equals_enkf.
keywords:
Ensemble Kalman Filter, Matheron Update, Gaussian Process, Data Assimilation, Geostatistics
@@ 1 Introduction
The Ensemble Kalman Filter (EnKF) (Evensen, 2003, 2009) is a cornerstone method that evolves ensembles of state vectors through model dynamics and updates them using observational data. The Matheron update provides a sample-based method for conditioning Gaussian random variables on observations (Doucet, 2010; Wilson et al., 2020, 2021), well-established in geostatistics but with little-known connections to ensemble data assimilation.
We establish that the ensemble update step in the EnKF is equivalent to an empirical Matheron update by putting them on a common probabilistic footing. This connection provides an alternative foundation for the EnKF and suggests improvements by leveraging computational optimizations from both communities.
1.1 Historical Context and Related Work
Although the affine residual update first appeared in geostatistics under the banner of conditioning by kriging (Chilès and Desassis, 2018), it was rediscovered many times:
  • Optimal interpolation (OI). Early numerical-weather-prediction (NWP) systems wrote the analysis as $x_{a}=x_{b}+K(y-Hx_{b})$ with a climatological Kalman gain; this is exactly the Matheron rule with fixed covariances (Hunt et al., 2007).
  • Stochastic EnKF. Evensen (2003) replaced the static climatological covariances of OI with an empirical ensemble, creating the now-ubiquitous EnKF.
  • Fast conditional simulation. Doucet (2010) noticed that one can obtain conditional Gaussian draws without a Cholesky factorisation by adding a linear residual—again the same formula.
  • Pathwise GP sampling. The machine-learning community re-invented the idea as path-wise conditioning for scalable Gaussian-process (GP) posterior draws (Wilson et al., 2020, 2021; Borovitskiy et al., 2023).
  • Reservoir inversion and EnRML. Iterative ensemble smoothers such as EnRML exploit the very same affine map in the context of PDE-constrained Bayesian inversion (Chen and Oliver, 2012).
  • Theory of ensemble inversion. Convergence proofs for ensemble inversion (Schillings and Stuart, 2017) and extensions via transport maps (Spantini et al., 2022) all start from the linear Matheron/EnKF identity.
This equivalence forms the algebraic core of at least six research communities, turning a patchwork of field-specific heuristics into transferable technology.
@@ 1.2 Notation
[…]
@@ 2 Matheron Update
[…]
@@
$\left(\boldsymbol{\mathsf{x}}\mid\boldsymbol{\mathsf{y}}{=}\boldsymbol{y}^{*}\right)$ $\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\boldsymbol{\mathsf{x}}+\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}\left(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}}\right).$ (2)
@@
$\operatorname{Var}\left[\boldsymbol{\mathsf{x}}+\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}})\right]$ $=\operatorname{Var}[\boldsymbol{\mathsf{x}}]+\operatorname{Var}[\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}})]$
$\hskip 20.00003pt+\operatorname{Var}(\boldsymbol{\mathsf{x}},\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}}))$
$\hskip 20.00003pt+\operatorname{Var}(\boldsymbol{\mathsf{x}},\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\mathrm{C}_{\boldsymbol{\mathsf{yy}}}^{-1}(\boldsymbol{y}^{*}-\boldsymbol{\mathsf{y}}))^{\top}$ (10)
@@
This pathwise approach to conditioning Gaussian variates has gained currency in machine learning as a tool for sampling and inference in Gaussian processes (Wilson et al., 2020, 2021), notably in generalising to challenging domains such as Riemannian manifolds (Borovitskiy et al., 2023).
@@ 3 Kalman Filter
We assume a known observation operator $\mathcal{H}$ such that observations $\boldsymbol{\mathsf{y}}$ are a priori related to state $\boldsymbol{\mathsf{x}}$ by $\boldsymbol{\mathsf{y}}=\mathcal{H}(\boldsymbol{\mathsf{x}})$; Thus there exists a joint law for the prior random vector
@@
[…]
@@
$\mathrm{K}$ $\coloneq\mathrm{C}_{\boldsymbol{\mathsf{xy}}}\,\mathrm{H}^{\top}\left(\mathrm{C}_{\boldsymbol{\mathsf{yy}}}\right)^{-1},$ (21)
$\coloneq\mathrm{C}_{\boldsymbol{\mathsf{xx}}}\,\mathrm{H}^{\top}\left(\mathrm{H}\,\mathrm{C}_{\boldsymbol{\mathsf{xx}}}\,\mathrm{H}^{\top}+\mathrm{R}\right)^{-1},$ (22)
is the Kalman gain. Hereafter we consider a constant diagonal $\mathrm{R}$ for simplicity, so that $\mathrm{R}=\rho^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$.
@@ 4 Ensemble Kalman Filter
$\bar{\mathrm{X}}\coloneq\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{x}^{(i)}\quad\quad\breve{\mathrm{X}}\coloneq\frac{1}{\sqrt{N-1}}\Bigl{(}\mathrm{X}-\bar{\mathrm{X}}\,\boldsymbol{1}^{\top}\Bigr{)}$ (24)
@@
$\widehat{\operatorname{Var}}(\boldsymbol{\mathsf{x}})\coloneq\operatorname{Var}_{\boldsymbol{\mathsf{x}}\sim\widehat{\mu}_{\mathrm{X}^{\prime}}}(\boldsymbol{\mathsf{x}})=\widehat{\mathrm{C}}_{\boldsymbol{\mathsf{xx}}}$ $\coloneq\frac{1}{N-1}\sum_{i=1}^{N}\left(\boldsymbol{x}^{(i)}-\widehat{\boldsymbol{m}}_{\boldsymbol{\mathsf{x}}}\right)\left(\boldsymbol{x}^{(i)}-\widehat{\boldsymbol{m}}_{\boldsymbol{\mathsf{x}}}\right)^{\top}$ (26)
@@
For posterior inference about $\boldsymbol{\mathsf{x}}$ given $\boldsymbol{\mathsf{y}}=\boldsymbol{y}^{*}$, we seek an ensemble $\mathrm{X}^{\prime}$ such that 111see Le Gland et al. (2011); Mandel et al. (2011); Kelly et al. (2014); Kwiatkowski and Mandel (2015); Del Moral et al. (2017) for finite-ensemble convergence
@@
$\mathrm{Y}\coloneq\mathcal{H}\mathrm{X}$ $\coloneq\begin{bmatrix}\mathcal{H}(\boldsymbol{x}^{(1)})&\mathcal{H}(\boldsymbol{x}^{(2)})&\dots&\mathcal{H}(\boldsymbol{x}^{(N)})\end{bmatrix}$. (31)
@@
[…]
@@
where $\rho^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$ is the observation error covariance matrix. The term $\upsilon^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$ comes from the regularization of the empirical observation covariance (as defined by the ensemble of $\mathrm{Y}$), while $\rho^{2}\mathrm{I}_{D_{\boldsymbol{\mathsf{y}}}}$ represents the known covariance of the observation noise. Combining these two gives the effective covariance in the observation space. [^2] The EnKF extends to nonlinear observation operators, though Gaussian assumptions no longer hold exactly (Evensen, 2009).
@@
5 Core result
Proposition 5.1.
The Empirical Matheron Update is equivalent to Ensemble Kalman Update
Proof 5.2.
@@ 6 Computational Complexity
The EnKF avoids computing full covariance matrices, instead using empirical ensemble statistics with cost $\mathcal{O}(D_{\boldsymbol{\mathsf{y}}}N^{2}+N^{3}+D_{\boldsymbol{\mathsf{x}}}D_{\boldsymbol{\mathsf{y}}}N)$. Since ensemble size $N$ is typically much smaller than dimensions $D_{\boldsymbol{\mathsf{x}}},D_{\boldsymbol{\mathsf{y}}}$, this is significantly more efficient than the naive Kalman update’s $\mathcal{O}(D_{\boldsymbol{\mathsf{y}}}^{3})$ cost.
7 Capabilities Unlocked by the Equivalence
(i) Linear-time GP sample paths.
EnKF algebra lets us compute the Matheron update with $\mathcal{O}(N^{2}d)$ cost—or $\mathcal{O}(Nd)$ with localisation—rather than $\mathcal{O}(d^{3})$ Cholesky factorizations, enabling multi-million-point GP draws (Wilson et al., 2020; Hunt et al., 2007).
(ii) Streaming and hyper-parameter learning.
Joint-state augmentation in EnKF permits on-line estimation of GP length-scales and noise variances (Kuzin et al., 2018), giving real-time Bayesian regression as data arrive.
(iii) Differentiable data-assimilation layers.
Because the affine map is a re-parameterisation, the complete update is compatible with auto-differentiation. One can embed an EnKF/Matheron block inside a neural network and back-propagate end-to-end, an approach already explored in differentiable simulators (Spantini et al., 2022; Schillings and Stuart, 2017).
(iv) Sparse covariance structure via localisation.
Local ensemble transform Kalman filters (LETKF) taper covariances in physical space, yielding block-banded precisions that transfer verbatim to spatial or graph-based GPs (Hunt et al., 2007).
(v) Rigorous error bounds.
Inverse-problem analyses prove contraction and bias properties of finite-ensemble EnKF (Schillings and Stuart, 2017); those proofs now bound the error of path-wise GP samplers for free.
Take-away.
Treating EnKF analysis as an empirical Matheron brings mature geophysical tool-kits—localisation, adaptive inflation, square-root variants—into scalable machine-learning while exporting automatic differentiation and sparse GP tricks back to data assimilation.
8 Numerical Illustration: 1D Kriging by Ensemble Kalman filer
We revisit the classic 1D kriging problem to demonstrate both the direct equivalence of the EnKF update to GP regression and the power of extending this link. We compare three solvers side-by-side: a standard GP implementation, its equivalent EnKF formulation, and the Local Ensemble Transform Kalman Filter (LETKF) (Bocquet et al., 2020; Hunt et al., 2007), a popular data assimilation variant that introduces covariance localization. This illustrates how mature geophysical tools can be immediately repurposed for scalable GP inference. Code is available in the supplement. We note that the Ensemble Kalman solvers used here are chosen for ease of implementation rather than efficiency of performance, and so we do not expect to see massive performance gains on these small problems.
Task. Recover a latent field $x\!\in\!\mathbb{R}^{d}$ on a unit grid from m=d/5m\!=\!d/5 noisy point samples (scaling with dimension) $y=Hx+\eta$, $\eta\!\sim\!\mathcal{N}(0,\tau^{2}I)$ with $\tau=0.2$. The prior is $\mathcal{N}(0,K)$, $K_{ij}=\sigma^{2}\!\exp(-\lVert r_{i}-r_{j}\rVert^{2}/2\ell^{2})$ with $\sigma=1$, $\ell=0.2$. We test grid sizes $d\in\{200,400,600,800\}$.
Methods.
  • GP Regression: The exact posterior computed via GaussianProcessRegressor from scikit-learn. This serves as our ground truth and scales as $\mathcal{O}(m^{3})$ with observations.
  • EnKF: An implementation of the standard Ensemble Kalman Filter using DAPPER (Raanes et al., 2024). It is the direct, empirical equivalent of the Matheron update and is expected to be computationally cheaper than the exact GP.
  • LETKF: The Local Ensemble Transform Kalman Filter (Hunt et al., 2007; Bocquet et al., 2020), also implemented in DAPPER. This method introduces localization, where each state variable is updated using only a subset of nearby observations. This is a common technique in data assimilation for improving accuracy in high-dimensional systems.
We report median fit and predict times obtained with our in-process timer over multiple runs to ensure statistical robustness.
Accuracy.
[…]
fig:posterior-samples shows posterior draws from all three methods using squared-exponential basis functions matched to the generating process (the “true prior” setting). We see similar RMSE for the mean estimate between the methods, with a slightly overconfident posterior uncertainty in the ensemble methods.
Speed.
Figure LABEL:fig:timing-curves shows separate training and prediction times. Training with scikit-learn scales cubically in $m$; prediction scales quadratically, but still rises steeply for large $m$. In contrast, EnKF and LETKF training scale linearly in ensemble size and prediction is effectively $O(1)$ for fixed ensemble, yielding two nearly flat curves. As the problem size scales beyond feasibility of GP regression, we expect to achieve efficiency gains by adopting mature DA techniques.
[…]
fig:posterior-samples Refer to caption
Figure 1: Posterior draws from the exact GP (sklearn), EnKF, and LETKF..
[…]
fig:timing-curvesRefer to caption
Figure 2: Fit (solid) vs. predict (dashed) wall times. Top: scaling with observations $m$. Bottom: scaling with state dimension $d$.
Take-away.
The results confirm that a standard EnKF is a computationally efficient method for path-wise GP sampling. More importantly, the equivalence acts as a bridge to a rich ecosystem of data assimilation techniques. The LETKF demonstrates this powerfully: by simply changing one line of code to invoke a localized filter, we gain access to a highly scalable inference method that parallels sparse or domain-decomposition approaches in the GP literature. This opens the door for the GP community to leverage decades of mature, practical DA engineering “for free.”